a x =3 y =c z ,b 2 =ac,y=

ax=3y=cz,b2=ac,y=

Grade:8

1 Answers

noogler
8 years ago
y=2xzlog(b)/(x+z)log(3)
 
let  ax=3y=cz=k(cnstant)
a=k1/x,3=k1/y,c=k1/z
given ac=b2
k1/x . k1/z=b2 
k1/x+1/z=b2
kx+z/xz=b2
k  =  b2xz/x+z
substitute k value in 3=k1/y
3=b2xz/(x+z)y
apply logarithm on both sides 
log3=2xzlog(b) /(x+z)y                                                                     log(ax)=xlog(a)
y=2xzlog(b)/(x+z)log(3)